OpenStudy (anonymous):

describe the vertical asymptote(s) and hole(s) for the graph: y=(x-4)/(x^2+3x+2)

OpenStudy (anonymous):

hint: y=(x-4)/(x^2+3x+2) = (x-4)/ [(x+1)(x+2)]

OpenStudy (anonymous):

I am honestly so lost, I have no idea what this means

OpenStudy (anonymous):

Oh, do you have to factor to get that answer?

OpenStudy (mathmale):

This is a situation in which your doing some online research is essential. You might want to Google "vertical asymptote" and learn that way what such an asymptote is and why we bother to find it (or them). Look at just what sourwing factored: the denominator (only). We are to find the values of x that make the denominator (only) = zero (0). Factoring the den. makes this easier. Take sourwing's two factors and set them equal to zero separately, and then solve for x in each case. Please type out your answers in the form x=a, x=b.

OpenStudy (anonymous):

x=-1 and x=-2

OpenStudy (anonymous):

But how do I figure out if, and where there are holes

OpenStudy (mathmale):

First: by typing out x=-1 and x=-2, you have correctly identified the two vertical asymptotes. Regarding "holes:" Look at the numerator of the given rational function you started with. What is the numerator?

OpenStudy (anonymous):

x-4

OpenStudy (mathmale):

Is x-4 equal to either of the two factors of the denominator?

OpenStudy (anonymous):

No, so is the hole -4, or 4?

OpenStudy (anonymous):

I don't really understand what a "hole" is, mathematically speaking

OpenStudy (mathmale):

"equal" = "the same as" Right. No. Therefore you will not be canceling any factors, right? You can't cancel x-4 and x=-1.

OpenStudy (mathmale):

Now suppose that your numerator were x-4 and the denom. were (x-4)(x+2). Could you cancel anything?

OpenStudy (anonymous):

the fours?

OpenStudy (anonymous):

because you would come up with a positive four in the denominator

OpenStudy (mathmale):

Not exactly, but you could cancel the FACTORS (x-4) and (x-4) in numerator and denom. respectively.

OpenStudy (mathmale):

No, you cancel only if the factors are exactly the same; You can cancel (x-4) / (x-4), but not (x-4) / (x+4).

OpenStudy (anonymous):

Oh, ok.

OpenStudy (mathmale):

So, in my example, can you, or can you not, cancel?

OpenStudy (anonymous):

you can't

OpenStudy (mathmale):

Let me write out my example again, for clarity:\[y=\frac{ x-4 }{ (x-4)(x+2) }\]

OpenStudy (mathmale):

Can you do any cancelling here?

OpenStudy (anonymous):

yes the x-4 and x-4 so you'd end up with just y=x+2

OpenStudy (mathmale):

In other words, y ou'd end up with \[y=\frac{ 1 }{ 1(x+2) }.\]

OpenStudy (mathmale):

Now, because the original function had (x-4) in both numerator and denominator, x could NOT be 4. Why not?

OpenStudy (anonymous):

Because you cancelled out the equation where x would equal 4, so it's not possible.

OpenStudy (mathmale):

Not quite. The reason is that if you have (x-4) in the denom., and then let x=4, you'll have zero in the denom. Right?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Not allowed. THAT is where your graph would have a hole!! You'd have a nice, smooth graph for x>-1, BUT still, you could not let x=4, because then the factor (x-4) in the original problem statement would be zero. Division by zero is not permitted. So, instead of drawing a nice, smooth graph right past x=4, you'd have to draw a hole in the graph where x=4.

OpenStudy (anonymous):

Oh! Okay, that makes sense

OpenStudy (mathmale):

So, in summary, your graph will have a hole in it if you have a factor in the denom. equal to a factor in the numerator. Even tho' you can cancel, you are not changing the fact that x can not equal 4 in the original function.

OpenStudy (mathmale):

u r awesome. Happy for the chance to work with you!

OpenStudy (anonymous):

Thank you!

OpenStudy (mathmale):

My pleasure! Hasta la vista!